# Elastohydrodynamics of 3D chemically active filaments

### v1

@@ 4 Discussion
In this work, we have investigated the dynamics that arise when coupling the hydrodynamic loads induced by diffusiophoretic motility with the elastic behaviour of flexible filaments. In order to explore of a large portion of the parameter space, we have opted for a coarse, rapid simulation methodology. This incorporates local versions of the slender body theories for hydrodynamic forcing and diffusion-dominated reactions, coupling resistive force theory with local Slender Phoretic Theory. From our validations of local Slender Phoretic Theory and previous validation of the approach to slender elastohydrodynamics, we expect that this provides good qualitative agreement and fair quantitative agreement with higher resolution methods. The key dimensionless quantity governing the strength of phoresis-induced hydrodynamic forces to elastic restoring forces, is the elastohydrodynamic number, $E_{h}$. By varying the key parameter of the system, the elastohydrodynamic number, $E_{h}$, across orders of magnitude, we have uncovered a profoundly rich array of dynamic behaviours that arise from the interplay between phoresis, hydrodynamics and three-dimensional elasticity.
@@
In this exploratory study, we focused on slender spheroidal filaments that naturally relax to straight configurations, and on two archetypical activity patterns: symmetric Saturn- and antisymmetric Janus-like catalytic coatings. In the rigid limit, these correspond to stationary pumps and self-propelling translators, respectively. When elasticity is introduced, these baseline behaviours evolve into a wide spectrum of non-linear dynamical states. In both examples, as the elastohydrodynamic number is increased, equivalent to the filaments becoming more active or more flexible, the dynamics of the filaments pass through several key dynamic regimes.
@@
At very low $E_{h}$, elasticity dominates activity and filaments relax towards their rigid behaviour. Once $E_{h}$ reaches a threshold value, we observe buckling phenomena that induce subsequent planar motion. For Janus filaments, this buckling precipitates the onset of swimming behaviour; for the Saturn filaments, this transitions the dynamics from translation to rotation. Increasing the elastohydrodynamic number further, early-time planar motion is seen to become unstable, and fully three-dimensional helical motion emerges. At very high $E_{h}$, the filaments are easily deformed by the activity-induced viscous forces, leading to rapid shape changes that appear as diffusive-like motion. This sequence of behaviours with increasing activity is similar to that seen in other models of active and forced elastic filaments, which often show regimes of linear, planar and three-dimensional motion (e.g.  Du Roure et al., 2019; Laskar et al., 2013; Clarke et al., 2024; Altunkeyik et al., 2025). However, we note that the details of the motion depend sensitively on the nature of the activity/forcing and the applied boundary conditions. For instance, in other systems we see different ordering of modes, at different parameter values, with different shape transitions.
@@
Although the general trends seen in both examples of activity profile were similar, including transitions from rigid to planar to helical and diffusive behaviours, there are some additional noteworthy dynamical regimes specific to each example at intermediate values of the elastohydrodynamic number. For example, there are a range of values of $E_{h}$, not far beyond the transition to three-dimensional motion, where we observed symmetric filaments settling into trapped periodic ‘pinwheeling’ orbits. These orbits emerge during the course of the motion, and appear to be stable attractors. We also observe a finite-width region of parameter space where the overall trajectories return to near-planar motion, with travelling waves moving from the filament centre to the tip, which induces a flapping-like motion from side-to-side. Similarly, planar motion is observed to return for the antisymmetric filaments, but only at the highest values of $E_{h}$ explored. However, the progression of shape changes during the motion in this case appears qualitatively different to the persistent motion of symmetric filaments. Further study is required to understand more fully why these intermediate dynamics emerge.
@@ Acknowledgements
Appendix A Verifying local SPT
A.1 Local vs non-local SPT
We verify the accuracy of using a local version of Slender Phoretic Theory by comparing the calculated slip flows with those computed by the non-local theory, which has been previously validated (Katsamba et al., 2020, 2024). An example of the calculated slip flows for each of our considered activity profiles are shown in Figure 13, for a fixed proscribed shape, showing good agreement between the local and non-local theories at the slenderness considered.
@@ A.1 Local vs non-local SPT
Figure 13: Comparison of leading-order slip flows calculated by local (solid) and non-local (dashed) SPT. (a) An S-shaped filament with $\mathcal{A}=\sqrt{1-s^{2}}$, (b) a helical filament with $\mathcal{A}=\sin{\pi s}$. In both cases, we used $N=40$ segments with $\epsilon=0.02$.
A.2 Convergence study

### v2

@@ 4 Discussion
In this work, we have investigated the dynamics that arise when coupling the hydrodynamic loads induced by diffusiophoretic motility with the elastic behaviour of flexible filaments. In order to explore a large portion of the parameter space, we have opted for a coarse, rapid simulation methodology. This incorporates local versions of the slender body theories for hydrodynamic forcing and diffusion-dominated reactions, coupling resistive force theory with local slender phoretic theory. From our validations of local slender phoretic theory and previous validation of the approach to slender elastohydrodynamics, we expect that this provides good qualitative agreement and fair quantitative agreement with higher resolution methods. The key dimensionless quantity governing the strength of phoresis-induced hydrodynamic forces to elastic restoring forces, is the elastohydrodynamic number, $E_{h}$. By varying this parameter across orders of magnitude, we have uncovered a profoundly rich array of dynamic behaviours that arise from the interplay between phoresis, hydrodynamics and three-dimensional elasticity.
@@
In this exploratory study, we focused on slender spheroidal filaments that naturally relax to straight configurations, and on two archetypical activity patterns: symmetric Saturn- and antisymmetric Janus-like catalytic coatings. In the rigid limit, these correspond to stationary pumps and self-propelling translators, respectively. When elasticity is introduced, these baseline behaviours evolve into a wide spectrum of non-linear dynamical states. In both examples, as the elastohydrodynamic number is increased, equivalent to the filaments becoming more active or more flexible, the dynamics of the filaments pass through several key dynamic regimes. These regimes are summarised in Figure 13.
@@
At very low $E_{h}$, elasticity dominates activity and filaments relax towards their rigid behaviour. Once $E_{h}$ reaches a threshold value, we observe buckling phenomena that induce subsequent planar motion. This buckling is reminiscent of classical Euler buckling in beams and rods as the compressive load (here, due to viscous forcing) is increased. For Janus filaments, this buckling precipitates the onset of swimming behaviour; for the Saturn filaments, this transitions the dynamics from translation to rotation. Increasing the elastohydrodynamic number further, early-time planar motion is seen to become unstable, and fully three-dimensional helical motion emerges. For the antisymmetric Saturn filaments, this occurred very soon after the buckling transition, and further study is needed to determine whether any planar states are stable on long-times for these filaments. We observe that the out-of-plane transition is typically preceded by a localised increase in torsion near the least curved section of the filament, and this instability may allow for the relaxation of elastic energies in highly curved filaments by deforming out-of plane. At very high $E_{h}$, the filaments are easily deformed by the activity-induced viscous forces, leading to rapid shape changes that appear as diffusive-like motion.
These types of behaviours are similar to those seen in other models of active and forced elastic filaments, which often show regimes of linear, planar and three-dimensional motion (e.g.  Du Roure et al., 2019; Laskar et al., 2013; Clarke et al., 2024; Altunkeyik et al., 2025). However, we note that the details of the motion depend sensitively on the nature of the activity/forcing and the applied boundary conditions. For instance, as the strength of activity is increased, clamped filaments with follower forces have been seen to exhibit an initial three-dimensional whirling motion before planar deformation (Clarke et al., 2024); instead we observe an initial planar instability, with three-dimensional motion occurring for higher values of activity. We expect that, in other systems, we shall see different ordering of modes, at different parameter values, with different shape transitions. Further study is needed to determine whether the behaviour seen here is characteristic of free-swimming elastic active filaments, or specific to chemoelastohydrodynamic filaments.
@@
Although the general trends seen in both examples of activity profile were similar, including transitions from rigid to planar to helical and diffusive behaviours, there are some additional noteworthy dynamical regimes specific to each example at intermediate values of the elastohydrodynamic number. For example, there are a range of values of $E_{h}$, not far beyond the transition to three-dimensional motion, where we observed symmetric filaments settling into trapped periodic ‘pinwheeling’ orbits. These orbits emerge during the course of the motion, and appear to be stable attractors. We also observe a finite-width region of parameter space where the overall trajectories return to near-planar ballistic motion, with travelling waves moving from the filament centre to the tip inducing a flapping-like motion from side-to-side. Further study is required to understand more fully why these dynamics emerge, along with the short or long-time persistence of apparent transients.
This study is a key step in understanding the behaviour of free-swimming chemically propelled elastic filaments. The results presented here are numerical experiments of an initial-value problem, and not a full analysis of the nonlinear dynamics. Hence, this study can be viewed as experimental evidence of the dynamics of these theoretical systems. Although we have made attempts to check the robustness of the presented behaviours to different conditions, it is possible that new dynamics may emerge, particularly on longer timescales. As such, we advocate for a detailed stability analysis of the dynamics, though this may be complicated by the non-small deviation of the dynamics from the elastic equilibrium. Such an analysis would likely be especially fruitful in the context of the antisymmetric activity profile considered in this work, where emergent behaviours were observed to depend on initial conditions.
Our numerics rely on asymptotically derived theories for both the chemical concentration and viscous flow around a thin and long filament. In each case, we use a local slender body theory, which leads to typical errors of a moderate size, $O(1/\log\epsilon)$, as discussed in appendix B. The errors are likely to be particularly acute when initially distant parts of the filaments become close together and non-local effects are more significant. In our simulations, this appears to occur when the filament is at its most deformable, such as the examples shown in fig. 12; this suggests that there may be higher errors associated with the results for extremely high values of $E_{h}$. We otherwise expect that our simulations provide reasonable accuracy for the dynamic behaviours observed across the wide range of parameter values considered.
@@ Acknowledgements
Appendix A Non-local SPT
For an axisymmetric activity, $\mathcal{A}(s)$, the first two terms of the surface concentration expansion for non-local slender phoretic theory, were previously calculated as (Katsamba et al., 2020)
${c}^{(0)}(s)$ $=\frac{1}{2}\rho(s)\mathcal{A}(s)\log\left(\frac{4}{\epsilon^{2}\rho^{2}(s)}\right)+\frac{1}{2}\int_{-1}^{1}\left[\frac{\rho(\tilde{s})\mathcal{A}(\tilde{s})}{|\bm{R}_{0}(s,\tilde{s})|}-\frac{\rho(s)\mathcal{A}(s)}{|s-\tilde{s}|}\right]\mathrm{d}\tilde{s}$ (20)
${c}^{(1)}(s,\theta)$ $=\frac{1}{2}\rho^{2}(s)\kappa(s)\mathcal{A}(s)\cos\Theta(s,\theta)\left[\log\left(\frac{4}{\epsilon^{2}\rho^{2}(s)}\right)-3\right]$
$\qquad-\rho(s)\int\limits_{-1}^{1}\!\left[\frac{\rho(\tilde{s})\mathcal{A}(\tilde{s})}{|\bm{R}_{0}(s,\tilde{s})|^{3}}\bm{R}_{0}(s,\tilde{s})\cdot\hat{\bm{e}}_{\rho}(s,\theta)+\frac{\rho(s)\kappa(s)\mathcal{A}(s)\cos\Theta(s,\theta)}{2|s-\tilde{s}|}\right]\mathrm{d}\tilde{s}$, (21)
where $\hat{\bm{e}}_{\rho}(s,\theta)$ is the unit vector pointing radially outwards from any cross-section. The local version of slender phoretic theory, used in this work, neglects the line integral contributions as they are size $\log(1/\epsilon)$ smaller than the logarithmic terms.
Appendix B Verifying local SPT
B.1 Local vs non-local SPT
We verify the accuracy of using a local version of slender phoretic theory by comparing the calculated slip flows with those computed by the non-local theory, which has been previously validated (Katsamba et al., 2020, 2024). An example of the calculated slip flows for each of our considered activity profiles are shown in Figure 14, for a fixed prescribed shape, showing good agreement between the local and non-local theories at the slenderness considered.
@@ B.1 Local vs non-local SPT
Figure 14: Comparison of leading order slip flows calculated by local (solid) and non-local (dashed) SPT. (a) An S-shaped filament with $\mathcal{A}=\sqrt{1-s^{2}}$, (b) a helical filament with $\mathcal{A}=\sin{\pi s}$. In both cases, we used a mobility $\mathcal{M}=+1$, with $N=40$ segments and slenderness $\epsilon=0.02$.
B.2 Convergence study